### v1

# QUEST#4X: an extension of QUEST#4 for benchmarking multireference wavefunction methods

Abstract
Given a number of datasets for evaluating the performance of single reference methods for the low-lying excited states of closed-shell molecules, a comprehensive dataset for assessing the performance of multireference methods for the low-lying excited states of open-shell systems is still lacking. For this reason, we propose an extension (QUEST#4X) of the radial subset of QUEST#4 [J. Chem. Theory Comput. 2020, 16, 3720] to cover 110 doublet and 39 quartet excited states. Near-exact results obtained by iCIPT2 (iterative configuration interaction with selection and second-order perturbation correction) are taken as benchmark to calibrate SDSCI (static-dynamic-static configuration interaction) and SDSPT2 (static-dynamic-static second-order perturbation theory), which are minimal MRCI and CI-like perturbation theory, respectively. It is found that SDSCI is very close in accuracy to ic-MRCISD (internally contracted multireference configuration interaction with singles and doubles), although its computational cost is just that of one iteration of the latter. Unlike most variants of MRPT2, SDSPT2 treats single and multiple states in the same way, and performs similarly as MS-NEVPT2 (multi-state n-electron valence second-order perturbation theory). These findings put the SDS family of methods (SDSPT2, SDSCI, and iCIPT2, etc.) on a firm basis.
1 Introduction
The last decades have witnessed fast progresses in the development of both wavefunction- and density-based quantum chemical methods for describing electronic structures of chemical systems. It is generally true that the strength and weakness of each of such methods should be uncovered before it can be applied safely to unknown problems. To this end, some standardized datasets should be established, such that they can be employed to identify the error bars of a given method. A dataset is composed of two ingredients, target systems and reference data. The former refers to chosen molecules and their properties, whereas the latter refers to corresponding experimental or highly accurate theoretical values. The target systems can be classified according to the simple criteria, (1) closed-shell (CS) or open-shell (OS), (2) main group (MG) or transition metal (TM), (3) ground state (GS) or excited state (ES), thereby leading to 8 types of datasets (see Fig. 1), to which the available datasets can be assigned. It can be seen from Table 1 that there exist 123 datasets for ground states of CS-MG 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101 , OS-MG 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 102 , and CS/OS-TM 1, 2, 3, 4, 5, 6, 7, 8, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123 (including lanthanides109, 118, 117 and actinides110, 116). Such datasets were mainly used to calibrate SR methods 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139 , especially density functional theory (DFT) 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82 , although some of them were also used to access the performance of MR methods 118, 117, 140, 115, 100, 141, 101, 142, 23, 22. In parallel, there exist 73 datasets oriented to excited states of CS-MG 4, 5, 6, 7, 8, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197 , OS-MG 4, 5, 6, 7, 8, 143, 144, 145, 198, 146, 147, 199, 200, 201, 202 , CS-TM 4, 5, 6, 7, 8, 143, 203, 204, 205, 206, 207 , and OS-TM 4, 5, 6, 7, 8, 143, 203, 204, 208, 209, 210 , aiming to calibrate both SR 4, 5, 6, 7, 8, 143, 144, 145, 146, 203, 204, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 198, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 199, 200, 201, 202, 205, 206, 207, 208, 209, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 215, 231, 232, 233, 234, 235, 236, 237, 238, 239 and MR 146, 147, 203, 204, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 205, 206, 207, 209, 210, 238, 240, 241, 242, 243, 244, 245, 246, 247, 248, 249, 250, 251, 252 methods.
[] Up-to-date datasets* Ground State CS-MG Refs. 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101 OS-MG Refs. 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 102 CS-TM Refs. 1, 2, 3, 4, 5, 6, 7, 8, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122 OS-TM Refs. 1, 2, 3, 4, 5, 6, 7, 8, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 123 Excited State CS-MG Refs. 4, 5, 6, 7, 8, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197 OS-MG Refs. 4, 5, 6, 7, 8, 143, 144, 145, 198, 146, 147, 199, 200, 201, 202 CS-TM Refs. 4, 5, 6, 7, 8, 143, 203, 204, 205, 206, 207 OS-TM Refs. 4, 5, 6, 7, 8, 143, 203, 204, 208, 209, 210
* CS: closed-shell; OS: open-shell; MG: main group; TM: transition metal.
As for the reference data, the very first point to be realized lies in that experimental measurements often cannot be taken as they stand. For instance, experimentally measured 0-0 transition energies cannot directly be compared to theoretically calculated vertical excitation energies (VEE), which are not observable. Instead, geometric and vibrational effects must be added to theoretical VEEs or subtracted from experimental 0-0 transition energies 253, 143, 171, 153, 156, 179, 166, 254, 176. The lack of sufficient experimental data necessitates the use of some high-level methods as the reference for low-level methods. However, this is somewhat empirical, if not arbitrary. In particular, it is often the case that the chosen high-level method is itself not sufficiently accurate. For instance, the near-exact iCIPT2/TZVP (iterative configuration interaction with selection and second-order perturbation255, 256) calculations147 of the singlet VEEs of the dataset196 overturned the recommendation of using the CASPT2/TZVP rather than the CC3/TZVP values as the theoretical best estimates (TBE),196 which was based simply on the extent of MR characters.
Close inspections of the up-to-date datesets (see Table 1) reveal that there exists only one dataset147 for calibrating MR methods for the excited states of open-shell organic systems. As such, more comprehensive datasets are highly desired. As a first try towards this goal, we start with the QUEST#4 dataset145, which contains 24 organic radicals (see Fig. S1 in the Supporting Information). Since only 51 doublet excited states were reported therein, an immediate extension is to cover more excited states, both doublet and quartet. Such extended dataset, covering 110 doublet and 39 quartet states for the 24 radicals, is to be dubbed QUEST#4X. The VEEs calculated by iCIPT2 will be employed to calibrate MR methods [SS/MS-NEVPT2 (single/muti-state second-order n-electron valence state perturbation theory)257, 258, SDSPT2 (static-dynamic-static (SDS) second-order perturbation theory)259, 260, SDSCI259, and ic-MRCISD (internally contracted multireference configuration interaction with singles and doubles)261], as well as the spin-adapted open-shell TD-DFT262, 263, 264. The SDSCI, SDSPT2, and iCIPT2 methods are first recapitulated in Sec. 2, which is followed by computational details in Sec. 3. The results are analyzed in Sec. 4. The paper is closed with concluding remarks in Sec. 5.
2 Computational Methods
The (restricted) static-dynamic-static (SDS) family of methods for strongly correlated systems of electrons have been detailed before147, 265. Therefore, it is only necessary to explain here their international relations. In the first place, no matter how many electrons and how many orbitals are to be correlated, SDSCI259 always constructs and diagonalizes a -by- matrix for states simultaneously. Therefore, SDSCI is a minimal MRCI resulting from the combination of the good of perturbation theory, intermediate Hamiltonian, and configuration interaction. Taking the eigenvectors of SDSCI as new references and repeating SDSCI until convergence, we obtain iCI266, where each iteration accesses a space that is higher by two ranks than that of the preceding iteration. That is, up to -tuple excitations (relative to the initial reference space) can be accessed if iterations are carried out. Because of the variational nature, any minor loss of accuracy stemming from the contractions can be removed by carrying out some micro-iterations. In other words, by controlling the numbers of macro- and micro-iterations, iCI will generate a series of contracted/uncontracted single/multireference CISD, with the resulting energy being physically meaningful at each level. It has been shown both theoretically and numerically that iCI can converge monotonically and quickly from above to FCI, even when starting with a very poor initial guess. As such, iCI can be interpreted as an exact solver of FCI. Further combined with the selection of configurations for static correlation and perturbation correction for dynamic correlation, we obtain the near-exact iCIPT2255, 256. On the other hand, if the block of the SDSCI matrix (i.e., matrix elements within the first-order interacting space) is replaced by , we obtain SDSPT2259, 260, a CI-like MRPT2 that treats single and multiple states in the same way and is particularly advantageous when a number of states are nearly degenerate (because of the sufficient relaxation of the reference coefficients)260. In short, if the SDSCI calculation is to be performed, we would obtain SDSPT2 results for free. The latter further yields, e.g., MS-NEVPT2 results for free, because all matrix elements required by MS-NEVPT2258 are already available. Moreover, SDSCI can be taken as a start of ic-MRCISD, so as to facilitate the convergence of the latter. As a matter of fact, the increments of the ic-MRCISD iterations clearly indicate the accuracy of SDSCI. Given so many good features, neither SDSCI nor SDSPT2 is size consistent. However, the errors can readily be cured147 by using the Pople correction267. This is also the case for ic-MRCISD261. Therefore, the SDSPT2, SDSCI, and ic-MRCISD results reported here all refer to those with the Pople corrections.
3 Computational Details
The 24 radicals and their geometries in the QUEST#4 dataset145 were held unchanged in the iCIPT2255, 256, ic-MRCISD261, SDSCI259, 260, SDSPT2260, and (MS-)NEVPT2257, 258 calculations carried out with the BDF program package268, 269, 270, 271, 272 under the highest Abelian group symmetries. Since the aug-cc-pVTZ (AVTZ) basis sets273, 274, 275 are good enough for most of the excitation energies (with the mean deviation (MD) only of 0.02 eV from the complete basis set (CBS) limit)145, they were also used here. SA-CASSCF (state-averaged complete active space self-consistent field) calculations with equal weights for all states were first carried out (see Sec.4.3 for the various active spaces). The Dyall Hamiltonian276 was then diagonalized within the chosen active space and used as the active part of in both SDSPT2 and (MS-)NEVPT2. The inactive part () of is composed of orbital energies for the doubly and zero occupied orbitals. They were obtained by diagonalizing the generalized Fock matrix (constructed with the state-averaged one-particle density matrix (1RDM)) within the doubly and zero occupied subspaces separately. However, such a choice of is problematic for pure Rydberg states when they were averaged equally with valence states in the SA-CASSCF calculations. The reason is very simple: pure Rydberg states stem from configurations very different from those of valence states. As such, it should be better to use state-specific orbital energies, that is, the doubly and zero occupied orbital energies are to be determined by the Fock matrix that is constructed with the 1RDM of each Rydberg state itself. Conceptually, the resulting state-specific orbitals should also be used when constructing the SDSPT2/MS-NEVPT2 effective Hamiltonian matrix. However, this is not only expensive due to additional integral transformations but also means that the off-diagonal Hamiltonian matrix elements would involve two sets of non-orthogonal orbitals. Therefore, SA-CASSCF instead of state-specific orbitals were still used in such ‘state-dependent Fock’ for 242. Although it is essentially indistinguishable from the ‘sate-averaged Fock’ for valence excitations242, it is a must for pure Rydberg states (see Table S1 in the Supporting Information).
Since iCIPT2 samples the whole Hilbert space, the initial SA-CASSCF calculation is not really needed, although it does provide a good start for the selection procedure. Anyway, iCIPT2 can work with natural orbitals (NOs) obtained by diagonalizing the 1RDM for the selected configurations with, e.g., . Here, is the single parameter involved in iCIPT2. That is, all configuration state functions with coefficients smaller in absolute value than will be pruned away from the variational space determined iteratively (for more details, see Ref. 256). With the fixed NOs, a series of iCIPT2 calculations can be performed with decreasing and are then extrapolated to the zero limit. The extrapolation is particularly robust when a linear fit is sufficient. It turns out that should be used for C3H5, CNO, CON, and CH2NO2, whereas are already sufficient for other systems.
It should also be mentioned that the core orbitals were kept frozen in the correlation calculations.
Finally, calculations with the spin-adapted open-shell TD-DFT262, 263, 264 under the Tamm-Dancoff approximation (denoted as X-TDA), in conjunction with the BHandHLYP functional 277, 278, 279 (the simplest yet reliable functional for the low-lying excited states of open-shell systems200, 201), were also performed for comparison.
4 Results and Discussion
4.1 Reference data
Ideally, well-resolved experimental data can be taken as the reference. However, for the 24 radicals considered here, only some experimental 0-0 transitions280, 281, 282, 283, 284, 285, 286, 287, 288, 289, 290, 291, after correcting the geometric and vibrational effects179, 202, 283, are available for comparison (see Table 4.1). Nevertheless, they are enough to verify the accuracy of iCIPT2. It can be seen from Table 4.1 that iCIPT2, in conjunction with the AVTZ basis sets273, 274, 275, is indeed very accurate, except for the state of CH3 and the state of NO. It turns out that there exist substantial basis set effects for the two states145. When the AV5Z basis sets273, 274, 275 are used, the errors for the two states are reduced from 0.21 and 0.25 eV to 0.04 and 0.06 eV, respectively. Not surprisingly, iCIPT2 is very close to CIPSI (configuration interaction using a perturbative selection made iteratively)292, since both are near-exact methods. The deviations of the methods (UCC3, UCCSDT, UCCSDTQ, and CIPSI) adopted in Ref. 145 from iCIPT2 are further plotted in Fig. 2 for the 51 doublets reported therein. All these pinpoint that the iCIPT2/AVTZ VEEs for the 110 doublets and 39 quartets of the 24 radicals can be taken as a solid reference for calibrating other methods with the same basis sets.
[] Experimental and theoretical vertical excitation energies (VEEs in eV) of 26 doublet states Radical State Expt. iCIPT2/AVTZa CCSDT/AVTZb CIPSI/AVTZb CIPSI/CBSc C3H5 3.41 d 3.42 3.43 - - BeF 6.16 e 6.27 6.23 6.21 - 4.14 e 4.15 4.15 4.14 4.13 BeH 2.48 e 2.49 2.49 2.49 2.48 6.32 e 6.46 6.45 6.46 6.46 CH 3.94 e 3.98 4.03 3.98 3.96 2.88 e 2.91 2.94 2.91 2.90 3.26 f 3.29 3.31 3.29 3.28 CH3 5.73 e 5.86 (5.88) 5.86 5.85 5.88 7.44 e 7.65 (7.48) 7.65 7.65 7.48 CN 3.22 e 3.22 3.25 3.22 3.21 1.32 e 1.34 1.38 1.34 1.33 CNO 1.55 g 1.62 1.71 1.61 1.61 CO+ 5.82 e 5.80 5.70 5.81 5.80 3.26 e 3.27 3.26 3.28 3.26 F2BO 2.77 h 2.77 2.78 - - 0.70 h 0.70 0.71 - - F2BS 2.91 i 2.94 2.93 - - 0.46 i 0.49 0.48 - - H2PS 2.77 j 2.73 2.75 2.72 2.71 NCO 2.80 k 2.87 2.87 2.83 2.89 NO 5.92 e 6.11 (6.05) 6.13 6.13 6.12 7.03 e 7.28 (7.09) 7.29 - [7.21] OH 4.09 e 4.11 4.12 4.10 4.09 PH2 2.74 l 2.76 2.77 2.77 2.76 C2H3 3.22 m 3.28 3.31 3.26 - MDn 0.05 0.06 0.05 0.04 MADn 0.06 0.06 0.06 0.06 SDn 0.09 0.10 0.08 0.08 MAXn 0.25 0.26 0.21 0.18
a
This work with AVTZ basis sets. In parentheses are AV5Z results. bRef. 145. cRef. 145. In square brackets is the CCSDTQ/AVQZ result. dRefs. 179, 280. eRefs. 202, 281, 282. fRefs. 283, 284. gRefs. 179, 285. hRefs. 179, 286. iRefs. 179, 287. jRefs. 179, 288. kRefs. 179, 289. lRefs. 179, 290. mRefs. 179, 291. nMD: mean deviation; MAD: mean absolute deviation; SD: standard deviation; MAX: maximum deviation. Unavailable data were excluded from the analysis.
4.2 QUEST#4X
Some criterion should first be established for adding additional excited states to the QUEST#4 dataset145. Since the highest VEE reported therein is 8.02 eV ( of OH), only those singly excited doublet and quartet states with VEEs up to 8.0 eV are to considered. To identify such states, iCIPT2/AVTZ calculations were performed on the low-lying excited states of the 24 radicals for each irreducible representation (irrep). Eventually, 59 doublet states (including 17 Rydberg states) and 39 quartet states (including 6 Rydberg states) were added to QUEST#4 (which has 51 doublet states, including 11 Rydberg states), leading to QUEST#4X. Although the AVTZ basis sets may not be sufficient for some of the Rydberg states, the CIPT2/AVTZ results can certainly be taken as reference for calibrating other methods with the same basis sets.
4.3 Active space
Unlike iCIPT2, the NEVPT2, SDSPT2, SDSCI, and ic-MRCISD calculations all require a proper choice of active spaces. The occupation numbers (ON) of the natural orbitals (NO) resulting from the iCIPT2 calculations were taken here as a guidance. Those NOs with ONs significantly different from zero or two were taken as the active orbitals, the number of which was, however, not allowed to exceed 14, in order to make the MR calculations feasible. The resulting active spaces for the 24 radicals are documented in Table 4.3. The SA-CASSCF calculations can be initiated and supervised by the iCAS approach (imposed automatic selection of complete active spaces)293, where once the valence atomic orbitals (see Table S2 in the Supporting Information) are prepared, all remaining steps are automated. In particular, the subspace matching algorithm therein preserves the characters of the orbitals (doubly occupied, active or virtual) throughout the self-consistent iterations.
[] Active spaces and electronic states by SA-CASSCF Radical Active Space Symmetry Irrep of Active MOa Electronic Stateb Doublet Quartet C3H5 (3e,9o) (2,2,1,4) (2,1,1,4) (0,1,1,0) BeF (9e,10o) (6,0,2,2) (2,0,1,1) - BeH (3e,11o) (5,0,3,3) (3,0,4,4) (0,0,1,1) BH2 (5e,10o) (5,0,3,2) (3,2,2,2) (0,1,0,0) CH (5e,9o) (5,0,2,2) (3,2,1,1) (0,1,1,1) CH3 (7e,11o) (6,0,3,2) (3,0,2,2) - CN (9e,8o) (4,0,2,2) (2,0,2,2) (2,2,0,0) CNO (15e,12o) (6,0,3,3) (3,2,3,3) (0,0,1,1) CON (15e,12o) (6,0,3,3) (1,1,2,2) (0,0,1,1) CO+ (9e,8o) (4,0,2,2) (2,0,1,1) (1,0,0,0) F2BO (19e,14o) (6,1,4,3) (2,1,2,1) - F2BS (19e,14o) (6,1,4,3) (1,0,2,1) (0,0,1,0) H2BO (11e,10o) (5,0,3,2) (2,0,3,2) (0,0,1,0) HCO (11e,10o) (8,2) (3,3) (1,1) HOC (11e,10o) (8,2) (3,3) (1,1) H2PO (13e,10o) (7,3) (4,3) (2,1) H2PS (13e,10o) (7,3) (3,2) (3,3) NCO (15e,12o) (6,0,3,3) (2,0,3,3) (0,0,1,1) NH2 (7e,10o) (5,0,3,2) (2,0,1,2) (0,0,0,1) CH2NO2 (13e,10o) (3,1,3,3) (2,2,2,2) (1,1,1,0) NO (11e,10o) (6,0,2,2) (2,0,2,2) (0,1,1,1) OH (7e,9o) (5,0,2,2) (1,1,1,1) (0,1,0,0) PH2 (7e,10o) (5,0,3,2) (2,1,2,2) (0,1,0,1) C2H3 (11e,12o) (10,2) (3,3) (1,1)
a
: (); : (). b: (); : ().
4.4 Vertical Excitation Energies
The SS-NEVPT2, MS-NEVPT2, SDSPT2, SDSCI, and ic-MRCISD calculations on the VEEs of the 110 doublet and 39 quartet states of the 24 radicals in QUEST#4145 were performed with the above setup of active spaces. As a matter of fact, the SDSCI calculations yield SDSPT2 and SS/MS-NEVPT2 results for free. SDSCI also serves as the initial step of ic-MRCISD. The so-calculated VEEs are documented in Tables LABEL:table-ene. Apart from the overall error analysis shown in Table 4.4, the internal consistency of the methods is also analyzed in Table 4.4. Scatter plots for the iCIPT2 and X-TDA/SS-NEVPT2/MS-NEVPT2/SDSPT2/SDSCI/ic-MRCISD VEEs are further depicted in Fig. 3.
It can be seen from Fig. 3 that the three variants of MRPT2 (i.e., SS-NEVPT2, MS-NEVPT2, and SDSPT2) have a strong linear correlation with iCIPT2, with R2 larger than 0.995, which is companied by a mean absolute deviation (MAD) of ca. 0.1 eV. As can be seen from Table 4.4, the VEEs calculated by the three MRPT2 are closely aligned for more than 80% of the excited states, with the differences in between being less than 0.05 eV. However, SS-NEVPT2 fails occasionally. For instance, it fails to reproduce the ordering of the (6.83 eV) and (6.71 eV) states of H2PO. Despite some overestimates, both MS-NEVPT2 and SDSPT2 reproduce the correct ordering for the two states (MS-NEVPT2: 6.61 and 7.04 eV; SDSPT2: 6.64 and 7.05 eV), as compared to the iCIPT2 results (6.20 and 6.81 eV). A similar situation occurs also to ( and under ) and ( and under ) states of CH3. These represent examples where dynamic correlation revises significantly the coefficients of the reference functions, as shown in Table 4.4. Here, it deserves to be emphasized that, although MS-NEVPT2 and SDSPT2 are hardly distinguishable for the systems considered here (see Table 4.4), SDSPT2 does outperform MS-NEVPT2 for situations with multiple nearly degenerate states260.
[] Wavefunctions of excited states of H2PO and CH3 via CASSCF, MS-NEVPT2, and SDSPT2 H2PO Ground state: (10a’)1(3a”)2 (3a”)2 (11a’)0 (10a’)1 (4a”)1 72% 15% 18% 62% 3% 75% 87% 3% 7% 73% 84% 5% CH3 Ground state: (3a1)2(1b1)2(1b2)1 (1b2)1 to (5a1)0 (3a1)2 (1b2)1 96% 0% 0% 95% 20% 76% 75% 20% 42% 54% 54% 41% (1b2)1 to (2b1)0 (1b1)2 (1b2)1 96% 0% 0% 95% 20% 76% 75% 20% 48% 48% 48% 47%
Both SDSCI and ic-MRCISD have a perfect linear correlation with iCIPT2, with R2 being ca. 0.999, which is companied by a MAD of ca. 0.05 eV. SDSCI agrees with ic-MRCISD within 0.15 eV for all the excited states, given that the computational cost of SDSCI is merely that of one iteration of ic-MRCISD. More specifically, SDSCI agrees with ic-MRCISD within 0.05, 0.10, and 0.15 eV for 75%, 95%, and 100% of the excited states, respectively (cf. Table 4.4). In particular, this is not fortuitous, but holds also for other systems147. However, it deserves to be pointed out that both SDSCI and ic-MRCISD perform worse than the three MRPT2 for the state of nitromethyl (CH2NO2), see Table LABEL:table-ene.
Compared to the above MR methods, X-TDA/BHandHLYP has a weaker linear correlation with iCIPT2, as reflected by a lower R2 of ca. 0.927. Moreover, the orderings of several states of BH2, CH3, H2PO, H2PS, and C2H3 were not reproduced correctly, as marked in Table LABEL:table-ene. Nonetheless, the accuracy of X-TDA for doublet and quartet excited states of radicals is comparable to that of TD-DFT for singlet and triplet excited states of closed-shell systems, as already noticed before200, 201.
| Radical | State | X-TDA | SS-NEVPT2 | MS-NEVPT2 | SDSPT2 | SDSCI | ic-MRCISD | iCIPT2 |
| allyl | -0.24 | 0.19 | 0.19 | 0.16 | -0.15 | -0.06 | 4.94 | |
| (C3H5) | -0.28 | 0.24 | 0.24 | 0.21 | -0.11 | -0.03 | 5.46 | |
| -0.36 | 0.18 | 0.18 | 0.16 | -0.17 | -0.08 | 5.61 | ||
| 0.07 | 0.16 | 0.15 | 0.10 | -0.13 | -0.05 | 3.40 | ||
| -0.39 | 0.41 | 0.07 | 0.22 | 0.03 | 0.08 | 5.71 | ||
| -0.60 | 0.47 | 0.12 | 0.15 | 0.07 | 0.07 | 6.53 | ||
| -1.52 | -0.64 | 0.07 | -0.07 | -0.08 | -0.11 | 7.80 | ||
| -0.31 | 0.22 | 0.22 | 0.19 | -0.25 | -0.12 | 5.98 | ||
| -0.50 | 0.24 | 0.24 | 0.21 | -0.27 | -0.14 | 7.68 | ||
| BeF | -0.33 | 0.01 | 0.01 | 0.01 | -0.03 | -0.03 | 6.27 | |
| 0.05 | 0.02 | 0.02 | 0.04 | 0.06 | 0.03 | 4.15 | ||
| BeH | -0.37 | 0.01 | 0.01 | 0.01 | 0.01 | 0.00 | 5.52 | |
| 0.45 | 0.19 | 0.19 | 0.19 | 0.14 | 0.00 | 5.72 | ||
| 0.08 | 0.03 | 0.03 | 0.03 | 0.01 | 0.00 | 2.49 | ||
| -0.76 | 0.00 | 0.00 | 0.01 | 0.01 | 0.00 | 6.46 | ||
| 0.44 | 0.24 | 0.23 | 0.23 | 0.16 | 0.01 | 7.34 | ||
| 0.12 | 0.19 | 0.20 | 0.21 | 0.11 | -0.03 | 7.80 | ||
| 0.24 | 0.03 | 0.03 | 0.03 | -0.01 | 0.00 | 5.88 | ||
| BH2 | -0.24 | -0.03 | -0.02 | -0.02 | -0.01 | 0.01 | 5.72 | |
| -0.35 | -0.04 | -0.03 | -0.03 | -0.02 | 0.01 | 6.75 | ||
| 0.17 | -0.03 | -0.02 | -0.02 | 0.00 | 0.01 | 6.41 | ||
| 0.08 | 0.06 | 0.06 | 0.06 | 0.04 | -0.03 | 6.95 | ||
| 0.19* | 0.05 | 0.05 | 0.06 | 0.04 | 0.00 | 6.38 | ||
| -0.40* | -0.04 | -0.03 | -0.04 | -0.03 | 0.00 | 6.58 | ||
| 0.12 | 0.00 | 0.00 | 0.00 | 0.00 | 0.01 | 1.18 | ||
| -0.37 | -0.06 | -0.05 | -0.05 | -0.02 | 0.01 | 6.87 | ||
| 0.19 | -0.03 | -0.02 | -0.02 | 0.00 | 0.01 | 5.35 | ||
| CH | -1.05 | 0.06 | 0.05 | 0.05 | 0.02 | 0.00 | 3.98 | |
| -0.10 | -0.02 | -0.02 | -0.02 | -0.01 | -0.01 | 6.49 | ||
| 0.47 | 0.03 | 0.03 | 0.03 | 0.01 | -0.01 | 2.91 | ||
| 0.10 | 0.02 | 0.02 | 0.03 | 0.01 | -0.01 | 3.29 | ||
| 0.11 | -0.03 | -0.03 | -0.04 | -0.01 | 0.01 | 0.72 | ||
| -0.18 | 0.03 | 0.03 | 0.02 | 0.00 | 0.00 | 7.66 | ||
| CH3 | -0.23 | -0.02 | -0.01 | -0.02 | -0.02 | 0.01 | 5.86 | |
| 0.04* | 0.17* | 0.07 | 0.09 | 0.11 | 0.04 | 6.96 | ||
| -0.45 * | -0.14* | -0.01 | -0.05 | -0.05 | -0.03 | 7.19 | ||
| -0.26 | -0.06 | -0.03 | -0.03 | -0.01 | 0.00 | 7.65 | ||
| CN | 0.36 | 0.02 | 0.04 | 0.05 | 0.02 | -0.01 | 3.22 | |
| -0.46 | 0.04 | 0.05 | 0.05 | 0.00 | -0.01 | 1.34 | ||
| 0.84 | 0.11 | 0.12 | 0.14 | 0.07 | 0.01 | 7.88 | ||
| -0.30 | 0.04 | 0.05 | 0.05 | 0.02 | 0.02 | 6.03 | ||
| -0.70 | 0.07 | 0.08 | 0.09 | 0.03 | 0.00 | 7.12 | ||
| -0.74 | 0.13 | 0.14 | 0.15 | 0.05 | 0.01 | 7.88 | ||
| CNO | 1.49 | -0.08 | -0.08 | -0.08 | 0.05 | 0.01 | 1.62 | |
| 0.33 | 0.10 | 0.11 | 0.11 | 0.06 | 0.08 | 7.63 | ||
| 0.75 | 0.10 | 0.11 | 0.13 | 0.12 | 0.07 | 7.69 | ||
| -0.43 | -0.19 | -0.18 | -0.16 | -0.18 | -0.18 | 7.82 | ||
| 0.31 | 0.04 | 0.05 | 0.05 | 0.02 | 0.02 | 5.48 | ||
| 0.00 | 0.09 | 0.10 | 0.11 | 0.05 | 0.06 | 6.18 | ||
| 0.08 | 0.07 | 0.08 | 0.09 | 0.04 | 0.06 | 5.70 | ||
| CON | 2.42 | -0.04 | -0.04 | -0.04 | 0.03 | 0.00 | 3.80 | |
| -0.47 | 0.04 | 0.04 | 0.07 | 0.07 | 0.05 | 7.15 | ||
| -0.08 | 0.04 | 0.04 | 0.05 | 0.05 | 0.05 | 3.47 | ||
| -0.20 | -0.03 | -0.03 | -0.01 | 0.01 | 0.02 | 2.74 | ||
| CO+ | 0.83 | 0.05 | 0.05 | 0.07 | 0.01 | -0.01 | 5.80 | |
| 0.10 | 0.03 | 0.04 | 0.05 | -0.01 | 0.00 | 3.27 | ||
| -0.27 | 0.06 | 0.07 | 0.08 | 0.02 | 0.01 | 7.27 | ||
| F2BO | 0.15 | 0.04 | 0.04 | 0.04 | 0.01 | 0.01 | 2.77 | |
| 1.25 | 0.21 | 0.21 | 0.23 | 0.04 | 0.11 | 7.79 | ||
| 1.23 | 0.21 | 0.21 | 0.22 | -0.04 | 0.07 | 6.94 | ||
| 1.37 | 0.26 | 0.26 | 0.27 | -0.02 | 0.08 | 6.53 | ||
| 0.07 | -0.01 | -0.01 | -0.01 | -0.01 | 0.00 | 0.70 | ||
| F2BS | 0.09 | 0.10 | 0.10 | 0.10 | -0.01 | 0.04 | 2.94 | |
| 0.09 | 0.08 | 0.08 | 0.10 | 0.16 | 0.12 | 7.02 | ||
| -0.04 | 0.05 | 0.05 | 0.05 | 0.00 | -0.01 | 0.49 | ||
| -0.03 | -0.04 | -0.04 | -0.03 | 0.08 | 0.05 | 6.39 | ||
| H2BO | 0.02 | 0.06 | 0.06 | 0.07 | 0.01 | 0.00 | 3.51 | |
| 1.16 | 0.06 | 0.07 | 0.09 | 0.07 | 0.04 | 7.25 | ||
| 1.03 | 0.00 | 0.00 | 0.00 | 0.04 | 0.02 | 4.61 | ||
| 1.09 | 0.05 | 0.06 | 0.08 | 0.00 | 0.03 | 7.64 | ||
| -0.21 | 0.04 | 0.04 | 0.05 | -0.02 | 0.00 | 2.17 | ||
| 0.98 | 0.08 | 0.08 | 0.11 | 0.07 | 0.05 | 6.13 | ||
| 0.73 | 0.07 | 0.07 | 0.09 | 0.00 | 0.05 | 6.86 | ||
| HCO | 0.24 | 0.17 | 0.17 | 0.17 | -0.05 | 0.03 | 5.45 | |
| 0.07 | 0.17 | 0.18 | 0.18 | 0.07 | 0.08 | 6.15 | ||
| 0.17 | -0.01 | -0.02 | 0.00 | 0.04 | 0.02 | 2.09 | ||
| -0.18 | 0.11 | 0.11 | 0.11 | -0.02 | 0.01 | 7.07 | ||
| -0.28 | 0.08 | 0.10 | 0.11 | -0.01 | 0.00 | 7.73 | ||
| -0.21 | 0.05 | 0.05 | 0.05 | -0.02 | 0.04 | 6.75 | ||
| -0.21 | 0.04 | 0.05 | 0.05 | 0.00 | 0.05 | 6.37 | ||
| HOC | 0.36 | 0.14 | 0.14 | 0.13 | -0.04 | 0.02 | 3.78 | |
| 0.11 | 0.13 | 0.13 | 0.14 | 0.09 | 0.04 | 5.63 | ||
| 0.01 | -0.01 | -0.01 | 0.00 | 0.02 | 0.00 | 0.92 | ||
| 0.20 | 0.07 | 0.07 | 0.10 | 0.06 | 0.03 | 6.13 | ||
| 0.28 | 0.15 | 0.15 | 0.17 | 0.08 | 0.06 | 7.55 | ||
| 0.24 | 0.18 | 0.18 | 0.17 | -0.02 | 0.05 | 7.73 | ||
| 0.05 | -0.08 | -0.08 | -0.08 | 0.00 | 0.02 | 3.84 | ||
| H2PO | -0.05 | 0.18 | 0.24 | 0.25 | 0.08 | 0.04 | 4.21 | |
| 0.41 | 0.08 | 0.18 | 0.20 | 0.06 | 0.05 | 4.78 | ||
| -0.03 | 0.14 | 0.21 | 0.23 | 0.02 | 0.03 | 5.66 | ||
| -0.01 | 0.11 | 0.17 | 0.16 | -0.04 | 0.01 | 2.81 | ||
| 0.37 | 0.63* | 0.41 | 0.44 | 0.30 | 0.21 | 6.20 | ||
| 0.65 | -0.11* | 0.23 | 0.23 | -0.04 | 0.05 | 6.81 | ||
| 0.06 | 0.18 | 0.24 | 0.25 | 0.02 | 0.10 | 7.28 | ||
| 0.57 | 0.27 | 0.32 | 0.32 | 0.10 | 0.18 | 7.64 | ||
| 0.53 | 0.16 | 0.22 | 0.23 | 0.02 | 0.12 | 6.31 | ||
| H2PS | 0.13 | 0.09 | 0.11 | 0.11 | 0.04 | 0.02 | 2.73 | |
| 0.12 | 0.19 | 0.20 | 0.21 | 0.02 | 0.02 | 4.57 | ||
| -0.09 | 0.08 | 0.09 | 0.09 | -0.02 | 0.02 | 1.14 | ||
| 0.59 | 0.12 | 0.13 | 0.14 | 0.04 | 0.09 | 5.42 | ||
| 0.66 | 0.17 | 0.18 | 0.17 | 0.03 | 0.10 | 5.74 | ||
| 0.68* | 0.23 | 0.19 | 0.22 | 0.07 | 0.07 | 6.02 | ||
| -0.76* | 0.34 | 0.40 | 0.40 | 0.15 | 0.17 | 6.76 | ||
| 1.09* | 0.21 | 0.19 | 0.20 | 0.10 | 0.10 | 5.08 | ||
| -0.34* | 0.46 | 0.49 | 0.51 | 0.23 | 0.22 | 5.73 | ||
| -0.05 | 0.14 | 0.16 | 0.16 | -0.02 | 0.06 | 6.72 | ||
| NCO | 0.22 | -0.04 | -0.03 | -0.03 | 0.05 | 0.01 | 2.87 | |
| 1.22 | 0.07 | 0.08 | 0.08 | 0.03 | 0.03 | 7.03 | ||
| 0.88 | 0.03 | 0.05 | 0.05 | 0.02 | 0.02 | 4.72 | ||
| 0.04 | 0.10 | 0.11 | 0.12 | 0.07 | 0.05 | 7.43 | ||
| -0.12 | 0.06 | 0.07 | 0.08 | 0.04 | 0.06 | 6.79 | ||
| NH2 | 0.08 | 0.02 | 0.02 | 0.02 | 0.00 | 0.00 | 2.12 | |
| 0.04 | 0.06 | 0.06 | 0.05 | 0.00 | 0.00 | 7.69 | ||
| 0.01 | 0.06 | 0.06 | 0.07 | 0.01 | 0.01 | 6.48 | ||
| 0.02 | 0.03 | 0.03 | 0.02 | -0.02 | 0.00 | 7.75 | ||
| 0.04 | 0.04 | 0.04 | 0.03 | 0.00 | 0.01 | 7.29 | ||
| nitromethyl | 0.85 | -0.03 | 0.04 | 0.00 | 0.22 | 0.16 | 2.49 | |
| (CH2NO2) | 0.35 | -0.04 | 0.02 | 0.00 | 0.06 | 0.08 | 5.19 | |
| 1.13 | -0.05 | -0.02 | -0.04 | 0.10 | 0.08 | 2.30 | ||
| 0.63 | 0.02 | 0.12 | 0.14 | 0.10 | 0.14 | 6.71 | ||
| 1.04 | -0.01 | 0.05 | 0.03 | 0.15 | 0.13 | 1.98 | ||
| 0.35 | 0.00 | 0.06 | 0.04 | 0.01 | 0.05 | 4.64 | ||
| 0.23 | -0.04 | 0.09 | 0.14 | 0.44 | 0.32 | 5.30 | ||
| 0.05 | 0.00 | 0.07 | 0.05 | 0.07 | 0.09 | 4.94 | ||
| -0.50 | -0.03 | 0.03 | 0.00 | -0.03 | 0.01 | 4.28 | ||
| 0.10 | 0.01 | 0.08 | 0.05 | 0.01 | 0.06 | 4.45 | ||
| NO | 0.46 | 0.19 | 0.19 | 0.19 | -0.09 | -0.01 | 6.11 | |
| 0.28 | 0.17 | 0.17 | 0.16 | -0.11 | -0.02 | 7.28 | ||
| -0.45 | -0.04 | -0.04 | -0.04 | -0.15 | -0.13 | 7.79 | ||
| -0.21 | -0.04 | -0.04 | -0.02 | 0.05 | 0.02 | 6.36 | ||
| -0.45 | 0.06 | 0.06 | 0.05 | -0.03 | 0.02 | 6.78 | ||
| OH | 0.04 | 0.01 | 0.01 | 0.01 | 0.01 | 0.00 | 4.11 | |
| 0.09 | 0.01 | 0.01 | 0.01 | -0.01 | -0.01 | 8.03 | ||
| 0.09 | 0.02 | 0.02 | 0.02 | 0.00 | -0.01 | 7.50 | ||
| PH2 | 0.05 | 0.07 | 0.07 | 0.07 | 0.04 | -0.01 | 2.76 | |
| -0.15 | 0.05 | 0.05 | 0.04 | 0.00 | 0.02 | 5.90 | ||
| 0.24 | 0.21 | 0.21 | 0.22 | 0.10 | 0.02 | 7.10 | ||
| 0.08 | 0.10 | 0.10 | 0.10 | 0.03 | -0.02 | 5.17 | ||
| -0.07 | 0.08 | 0.09 | 0.09 | 0.02 | -0.02 | 5.75 | ||
| -0.14 | 0.06 | 0.06 | 0.06 | 0.01 | 0.01 | 7.47 | ||
| 0.09 | 0.04 | 0.04 | 0.04 | 0.00 | 0.03 | 6.16 | ||
| -0.03 | 0.13 | 0.13 | 0.13 | 0.06 | 0.05 | 6.95 | ||
| vinyl | 0.75* | 0.10 | 0.09 | 0.11 | 0.02 | 0.02 | 5.60 | |
| (C2H3) | -0.80* | 0.13 | 0.17 | 0.15 | -0.06 | 0.01 | 6.19 | |
| -0.10 | 0.04 | 0.05 | 0.05 | 0.01 | -0.03 | 3.28 | ||
| -0.03 | 0.05 | 0.06 | 0.09 | 0.08 | 0.02 | 4.70 | ||
| -0.42 | 0.14 | 0.16 | 0.15 | -0.07 | 0.02 | 7.54 | ||
| -0.15 | 0.05 | 0.06 | 0.05 | -0.03 | 0.02 | 4.54 | ||
| -0.38 | 0.16 | 0.17 | 0.15 | -0.08 | 0.03 | 7.40 |
Incorrect ordering.
[] Statistical analysis of the errors (in eV) of X-TDA, SS-NEVPT2, MS-NEVPT, SDSPT2, SDSCI, and ic-MRCISD relative to iCIPT2 for the vertical excitation energies of 110 doublet and 39 quartet states∗ Type X-TDA SS-NEVPT2 MS-NEVPT2 SDSPT2 SDSCI ic-MRCISD Doublets MD 0.17 0.07 0.08 0.08 0.02 0.02 (110 states) MAD 0.40 0.10 0.09 0.10 0.06 0.04 SD 0.58 0.15 0.12 0.12 0.08 0.06 MAX 2.42 0.63 0.41 0.44 0.44 0.32 Quartets MD -0.03 0.09 0.11 0.11 0.01 0.04 (39 states) MAD 0.31 0.11 0.12 0.12 0.05 0.06 SD 0.41 0.15 0.16 0.16 0.08 0.08 MAX 1.09 0.46 0.49 0.51 0.23 0.22 Overall MD 0.12 0.07 0.09 0.09 0.02 0.03 (149 states) MAD 0.38 0.10 0.10 0.10 0.05 0.04 SD 0.54 0.15 0.13 0.13 0.08 0.07 MAX 2.42 0.63 0.49 0.51 0.44 0.32
MD: mean deviation; MAD: mean absolute deviation; SD: standard deviation; MAX: maximum deviation.
[] Internal consistency (percentage number of states within a deviation range in absolute value) across SS-NEVPT2, MS-NEVPT2, SDSPT2, SDSCI, ic-MRCISD, and iCIPT2 Deviation Range (eV) SS-NEVPT2 vs MS-NEVPT2 SS-NEVPT2 vs SDSPT2 MS-NEVPT2 vs SDSPT2 MS-NEVPT2 vs SDSCI SDSPT2 vs SDSCI SDSCI vs ic-MRCISD ic-MRCISD vs iCIPT2 0.05 84 % 89 % 99 % 50 % 44 % 75 % 74% 0.10 95 % 95 % 99 % 72 % 68 % 95 % 89% 0.15 97 % 96 % 100 % 82 % 81 % 100 % 95%
5 Conclusions
A comprehensive dataset, QUEST#4X, has been introduced for the purpose of calibrating multireference (MR) methods. It encompasses 110 doublet and 39 quartet excited states for the 24 radicals in the dataset QUEST#4 (which has only 51 doublet states). The iCIPT2/AVTZ vertical excitation energies reported here serve as benchmark for any other MR methods. It is of interest to note that the MR methods considered here, i.e., NEVPT2, SDSPT2, SDSICI, and ic-MRCISD, share the same function space and differ only in the determination of the expansion coefficients. Because of this, SDSCI can be taken as the initial step of ic-MRCISD and can even yield SDSPT2 and MS-NEVPT2 results automatically. Moreover, SDSPT2 also yields MS-NEVPT2 results automatically. More interestingly, the degree of agreement between SDSPT2 and MS-NEVPT2, the magnitude of deviations of SDSPT2/MS-NEVPT2 from SDSCI, as well as the iterative increments going from SDSCI to ic-MRCISD are clear indicators for the internal consistency between the methods and hence their accuracy. A good internal consistency can only be achieved by choosing appropriate active spaces for all target states, as done here. Notwithstanding this, more challenging datasets for the excited states of open-shell transition metals and -elements remain to be established. Work along this direction is being carried out at our laboratory.