### v2

# QUEST#4X: an extension of QUEST#4 for benchmarking multireference wavefunction methods

Abstract
Given a number of datasets for evaluating the performance of single reference methods for the low-lying excited states of closed-shell molecules, a comprehensive dataset for assessing the performance of multireference methods for the low-lying excited states of open-shell systems is still desired. For this reason, we propose an extension (QUEST#4X) of the radical subset of QUEST#4 [J. Chem. Theory Comput. 2020, 16, 3720] to cover 110 doublet and 39 quartet excited states. Near-exact results obtained by iCIPT2 (iterative configuration interaction with selection and second-order perturbation correction) are taken as benchmark to calibrate SDSCI (static-dynamic-static configuration interaction) and SDSPT2 (static-dynamic-static second-order perturbation theory), which are minimal MRCI and CI-like perturbation theory, respectively. It is found that SDSCI is very close in accuracy to ic-MRCISD (internally contracted multireference configuration interaction with singles and doubles), although its computational cost is just that of one iteration of the latter. Unlike most variants of MRPT2, SDSPT2 treats single and multiple states in the same way, and performs similarly as MS-NEVPT2 (multi-state n-electron valence second-order perturbation theory). These findings put SDSCI and SDSPT2 on a firm basis.
1 Introduction
The last decades have witnessed fast progresses in the development of both wavefunction- and density-based quantum chemical methods for describing electronic structures of chemical systems. It is generally true that the strengths and weaknesses of each of such methods should be uncovered before they can be applied safely to unknown problems. To this end, some standardized datasets should be established, such that they can be employed to identify the error bars of a given method. A dataset is composed of two ingredients, target systems and reference data. The former refers to chosen molecules and their properties, whereas the latter refers to corresponding experimental or highly accurate theoretical values. The target systems can be classified according to the simple criteria, (1) closed-shell (CS) or open-shell (OS), (2) main group (MG) or transition metal (TM), (3) ground state (GS) or excited state (ES), thereby leading to 8 types of datasets (see Fig. 1), to which the available datasets can be assigned. It can be seen from Table 1 that there exist at least 123 datasets for ground states of CS-MG 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101 , OS-MG 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 102 , and CS/OS-TM 1, 2, 3, 4, 5, 6, 7, 8, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123 (including lanthanides109, 118, 117 and actinides110, 116). Such datasets were mainly used to calibrate SR methods 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139 , especially density functional theory (DFT) 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82 , although some of them were also used to assess the performance of MR methods 118, 117, 140, 115, 100, 141, 101, 142, 23, 22. In parallel, there exist 73 datasets oriented to excited states of CS-MG 4, 5, 6, 7, 8, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197 , OS-MG 4, 5, 6, 7, 8, 143, 144, 145, 198, 146, 147, 199, 200, 201, 202 , CS-TM 4, 5, 6, 7, 8, 143, 203, 204, 205, 206, 207 , and OS-TM 4, 5, 6, 7, 8, 143, 203, 204, 208, 209, 210 , aiming to calibrate both SR 4, 5, 6, 7, 8, 143, 144, 145, 146, 203, 204, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 198, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 199, 200, 201, 202, 205, 206, 207, 208, 209, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 215, 231, 232, 233, 234, 235, 236, 237, 238, 239 and MR 146, 147, 203, 204, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 205, 206, 207, 209, 210, 238, 240, 241, 242, 243, 244, 245, 246, 247, 248, 249, 250, 251, 252 methods.
[] Up-to-date datasets* Ground State CS-MG Refs. 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101 OS-MG Refs. 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 102 CS-TM Refs. 1, 2, 3, 4, 5, 6, 7, 8, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122 OS-TM Refs. 1, 2, 3, 4, 5, 6, 7, 8, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 123 Excited State CS-MG Refs. 4, 5, 6, 7, 8, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197 OS-MG Refs. 4, 5, 6, 7, 8, 143, 144, 145, 198, 146, 147, 199, 200, 201, 202 CS-TM Refs. 4, 5, 6, 7, 8, 143, 203, 204, 205, 206, 207 OS-TM Refs. 4, 5, 6, 7, 8, 143, 203, 204, 208, 209, 210
* CS: closed-shell; OS: open-shell; MG: main group; TM: transition metal.
As for the reference data, the very first point to be realized lies in that experimental measurements often cannot be taken as they stand. For instance, experimentally measured 0-0 transition energies cannot directly be compared to theoretically calculated vertical excitation energies (VEE), which are not observable. Instead, geometric and vibrational effects must be added to theoretical VEEs or subtracted from experimental 0-0 transition energies 253, 143, 171, 153, 156, 179, 166, 254, 176. The lack of sufficient experimental data necessitates the use of some high-level methods as the reference for low-level methods. However, this is not always reliable. It is often the case that the chosen high-level method is itself not sufficiently accurate. For instance, the near-exact iCIPT2/TZVP (iterative configuration interaction with selection and second-order perturbation255, 256) calculations147 of the singlet VEEs of the dataset196 overturned the recommendation of using the CASPT2/TZVP rather than the CC3/TZVP values as the theoretical best estimates (TBE),196 which was based simply on the extent of MR characters.
Close inspections of the up-to-date datasets (see Table 1) reveal that there exists only one dataset147 for calibrating MR methods for the excited states of open-shell organic systems. As such, more comprehensive datasets are highly desired. As a first try towards this goal, we start with the QUEST#4 dataset145, which contains 24 organic radicals (see Fig. S1 in the Supporting Information). Since only 51 doublet excited states were reported therein, an immediate extension is to cover more excited states, both doublet and quartet. Such an extended dataset, covering 110 doublet and 39 quartet states for the 24 radicals, is to be dubbed QUEST#4X. The VEEs calculated by iCIPT2 will be employed to calibrate MR methods [SS/MS-NEVPT2 (single/multi-state second-order n-electron valence state perturbation theory)257, 258, SDSPT2 (static-dynamic-static (SDS) second-order perturbation theory)259, 260, SDSCI259, and ic-MRCISD (internally contracted multireference configuration interaction with singles and doubles)261], as well as the spin-adapted open-shell TD-DFT262, 263, 264. The SDSCI, SDSPT2, and iCIPT2 methods are first recapitulated in Sec. 2, which is followed by computational details in Sec. 3. The results are analyzed in Sec. 4. The paper is closed with concluding remarks in Sec. 5.
2 Computational Methods
The (restricted) static-dynamic-static (SDS) family of methods for strongly correlated systems of electrons has been presented at length before147, 265. Therefore, it is only necessary to explain here their internal relations. In the first place, no matter how many electrons and how many orbitals are to be correlated, SDSCI259 always constructs and diagonalizes a -by- matrix for states simultaneously. Therefore, SDSCI is a minimal MRCI resulting from the combination of the good of perturbation theory, intermediate Hamiltonian, and configuration interaction. Taking the eigenvectors of SDSCI as new references and repeating SDSCI until convergence, we obtain iCI266, where each iteration accesses a space that is higher by two ranks than that of the preceding iteration. That is, up to -tuple excitations (relative to the initial reference space) can be accessed if iterations are carried out. Because of the variational nature, any minor loss of accuracy stemming from the contractions can be removed by carrying out some micro-iterations. In other words, by controlling the numbers of macro- and micro-iterations, iCI will generate a series of contracted/uncontracted single/multireference CISD, with the resulting energy being physically meaningful at each level. It has been shown both theoretically and numerically that iCI can converge monotonically and quickly from above to FCI, even when starting with a very poor initial guess. As such, iCI can be interpreted as an exact solver of FCI. Further combined with the selection of configurations for static correlation and perturbation correction for dynamic correlation, we obtain the near-exact iCIPT2255, 256. On the other hand, if the block of the SDSCI matrix (i.e., matrix elements within the first-order interacting space) is replaced by , we obtain SDSPT2259, 260, a CI-like MRPT2 that treats single and multiple states in the same way and is particularly advantageous when a number of states are nearly degenerate (because of the sufficient relaxation of the reference coefficients)260. In short, if the SDSCI calculation is to be performed, we would obtain SDSPT2 results for free. The latter further yields, e.g., MS-NEVPT2 results for free, because all matrix elements required by MS-NEVPT2258 are already available. Moreover, SDSCI can be taken as a start of ic-MRCISD, so as to facilitate the convergence of the latter. As a matter of fact, the increments of the ic-MRCISD iterations clearly indicate the accuracy of SDSCI. Given so many good features, neither SDSCI nor SDSPT2 is size consistent. However, the errors can readily be cured147 by using the Pople correction267. This is also the case for ic-MRCISD261. Therefore, the SDSPT2, SDSCI, and ic-MRCISD results reported here all refer to those with the Pople corrections.
3 Computational Details
The 24 radicals and their geometries in the QUEST#4 dataset145 were held unchanged in the iCIPT2255, 256, ic-MRCISD261, SDSCI259, 260, SDSPT2260, and (MS-)NEVPT2257, 258 calculations carried out with the BDF program package268, 269, 270, 271, 272 under the highest Abelian group symmetries. Since the aug-cc-pVTZ (AVTZ) basis sets273, 274, 275 are good enough for most of the excitation energies (with the mean deviation (MD) only of 0.02 eV from the complete basis set (CBS) limit)145, they were also used here. SA-CASSCF (state-averaged complete active space self-consistent field) calculations with equal weights for all states were first carried out (see Sec.4.3 for the various active spaces). The Dyall Hamiltonian276 was then diagonalized within the chosen active space and used as the active part of the zeroth-order Hamiltonian in both SDSPT2 and (MS-)NEVPT2. The inactive part () of is composed of orbital energies for the doubly and zero occupied orbitals. They were obtained by diagonalizing the generalized Fock matrix (constructed with the state-averaged one-particle density matrix (1RDM)) within the doubly and zero occupied subspaces separately. However, such a choice of is problematic for states dominated by Rydberg characters (more than 70%) when they were averaged equally with valence states in the SA-CASSCF calculations. The reason is very simple: Rydberg states stem from configurations very different from those of valence states. As such, it should be better to use state-specific orbital energies, that is, the doubly and zero occupied orbital energies are to be determined by the Fock matrix that is constructed with the 1RDM of each Rydberg state itself. Conceptually, the resulting state-specific orbitals should also be used when constructing the SDSPT2/MS-NEVPT2 effective Hamiltonian matrix. However, this is not only expensive due to additional integral transformations but also means that the off-diagonal Hamiltonian matrix elements would involve two sets of non-orthogonal orbitals. Therefore, SA-CASSCF instead of state-specific orbitals were still used in such ‘state-dependent Fock’ for 242. Although it is essentially indistinguishable from the ‘state-averaged Fock’ for valence excitations242 , it is a must for pure Rydberg states (see Table S1 in the Supporting Information). Nevertheless, the ‘state-averaged Fock’ still performs better for states of heavily mixed valence and Rydberg characters (e.g., the , , and states of C3H5 are roughly 1:1 mixtures of valence and Rydberg excitations).
Since iCIPT2 samples the whole Hilbert space, the initial SA-CASSCF calculation is not really needed, although it does provide a good start for the selection procedure. Anyway, iCIPT2 can work with natural orbitals (NOs) obtained by diagonalizing the 1RDM for the selected configurations with, e.g., . Here, is the single parameter involved in iCIPT2. That is, all configuration state functions with coefficients smaller in absolute value than will be pruned away from the variational space determined iteratively (for more details, see Ref. 256). With the fixed NOs, a series of iCIPT2 calculations can then be performed with decreasing , so as to approach the FCI limit () by linear fit of the iCIPT2 vs PT2 energy plot. For both the ground and excited states, the extrapolation uncertainties at the 95% confidence level were kept below 1 mHartree. Such uncertainties for the ground and excited states were then summed up as the error bars (less than 2 mHartree or 0.05 eV) for the extrapolated VEEs. The extrapolation distances, i.e., the differences between the largest iCIPT2 calculations (at ) and extrapolated values, are also provided for reference.
It should also be mentioned that the core orbitals were kept frozen in the correlation calculations.
Finally, calculations with the spin-adapted open-shell TD-DFT262, 263, 264 under the Tamm-Dancoff approximation (denoted as X-TDA), in conjunction with the BHandHLYP functional 277, 278, 279 (the simplest yet reliable functional for the low-lying doublet-doublet200 and doublet-quartet excitations201 of radicals), were also performed for comparison.
4 Results and Discussion
4.1 Reference data
Ideally, well-resolved experimental data can be taken as the reference. However, for the 24 radicals considered here, only some experimental 0-0 transitions280, 281, 282, 283, 284, 285, 286, 287, 288, 289, 290, 291, after correcting the geometric and vibrational effects179, 202, 283, are available for comparison (see Table 4.1). Nevertheless, they are enough to verify the accuracy of iCIPT2. It can be seen from Table 4.1 that iCIPT2, in conjunction with the AVTZ basis sets273, 274, 275, is indeed very accurate, except for the state of CH3 and the state of NO. It turns out that there exist substantial basis set effects for the two states145. When the AV5Z basis set273, 274, 275 is used, the errors for the two states are reduced from 0.21 and 0.25 eV to 0.04 and 0.06 eV, respectively. Not surprisingly, the iCIPT2 results are very close to those by selected CI (sCI)145 since both methods are nearly exact. The deviations of the methods (UCC3, UCCSDT, UCCSDTQ, and sCI) adopted in Ref. 145 from iCIPT2 are further plotted in Fig. 2 for the 51 doublets reported therein. All these pinpoint that the iCIPT2/AVTZ VEEs for the 110 doublets and 39 quartets of the 24 radicals can be taken as a solid reference for calibrating other methods with the same basis sets.
[] Experimental and theoretical vertical excitation energies (VEEs in eV) of 26 doublet states Radical State Expt. iCIPT2/AVTZa 95% intervalb CCSDT/AVTZc sCI/AVTZa,d TBE/CBSe C3H5 3.41 f 3.40 0.02 0.03 3.43 - - BeF 6.16 g 6.27 0.00h 0.01 6.23 6.21 0.02 - 4.14 g 4.15 0.00 0.00 4.15 4.14 0.01 4.13 BeH 2.48 g 2.49 0.00 0.00 2.49 2.49 0.00 2.48 6.32 g 6.46 0.00 0.00 6.45 6.46 0.00 6.46 CH 3.94 g 3.98 0.00 0.00 4.03 3.98 0.00 3.96 2.88 g 2.91 0.00 0.00 2.94 2.91 0.00 2.90 3.26 i 3.29 0.00 0.00 3.31 3.29 0.00 3.28 CH3 5.73 g 5.86 0.00 (5.88 0.00) 0.00 ( 0.01) 5.86 5.85 0.01 5.88 7.44 g 7.65 0.00 (7.48 0.00) 0.00 ( 0.01) 7.65 7.65 0.01 7.48 CN 3.22 g 3.22 0.00 0.00 3.25 3.22 0.00 3.21 1.32 g 1.34 0.00 0.00 1.38 1.34 0.01 1.33 CNO 1.55 j 1.62 0.01 0.00 1.71 1.61 0.01 1.61 CO+ 5.82 g 5.80 0.00 0.00 5.70 5.81 0.00 5.80 3.26 g 3.27 0.00 0.00 3.26 3.28 0.00 3.26 F2BO 2.77 k 2.77 0.01 0.04 2.78 - - 0.70 k 0.70 0.04 0.03 0.71 - - F2BS 2.91 l 2.95 0.02 0.05 2.93 - - 0.46 l 0.50 0.00 0.04 0.48 - - H2PS 2.77 m 2.73 0.01 0.01 2.75 2.72 0.02 2.71 NCO 2.80 n 2.87 0.01 0.03 2.87 2.83 0.05 2.89 NO 5.92 g 6.11 0.01 (6.05 0.01) 0.01 ( 0.04) 6.13 6.13 0.02 6.12 7.03 g 7.28 0.01 (7.09 0.01) 0.00 ( 0.04) 7.29 - [7.21] OH 4.09 g 4.11 0.00 0.00 4.12 4.10 0.01 4.09 PH2 2.74 o 2.76 0.00 0.00 2.77 2.77 0.00 2.76 C2H3 3.22 p 3.28 0.00 0.01 3.31 3.26 0.02 - MDq 0.05 0.06 0.05 0.04 MADq 0.06 0.06 0.06 0.06 SDq 0.09 0.10 0.08 0.08 MAXq 0.25 0.26 0.21 0.18
a
means extrapolation distance . In parentheses are the AV5Z results. bSum of the 95% confidence intervals of the ground and excited states. In parentheses are the AV5Z results. cRef. 145. dRef. 145. eRef. 145. sCI results calculated or corrected to at least AVQZ level. In brackets is the CCSDTQ/AVQZ result. fRefs. 179, 280. gRefs. 202, 281, 282. h6.22 0.00 eV with the AVTZ from Basis Set Exchange292. iRefs. 283, 284. jRefs. 179, 285. kRefs. 179, 286. lRefs. 179, 287. mRefs. 179, 288. nRefs. 179, 289. oRefs. 179, 290. pRefs. 179, 291. qMD: mean deviation; MAD: mean absolute deviation; SD: standard deviation; MAX: maximum deviation. Unavailable data were excluded from the analysis.
4.2 QUEST#4X
Some criteria should first be established for adding additional excited states to the QUEST#4 dataset145. Since the highest VEE reported therein is 8.02 eV ( of OH), only those singly excited doublet and quartet states with VEEs up to 8.0 eV are to be considered here. To identify such states, iCIPT2/AVTZ calculations were performed on the low-lying excited states of the 24 radicals for each irreducible representation (irrep). Eventually, 59 doublet states (including 17 Rydberg states) and 39 quartet states (including 6 Rydberg states) were added to QUEST#4 (which has 51 doublet states, including 11 Rydberg states), leading to QUEST#4X. Although the AVTZ basis sets may not be sufficient for some of the Rydberg states, the iCIPT2/AVTZ results can certainly be taken as the reference for calibrating other methods with the same basis sets.
4.3 Active space
Unlike iCIPT2, the NEVPT2, SDSPT2, SDSCI, and ic-MRCISD calculations all require a proper choice of active spaces. The occupation numbers (ON) of the natural orbitals (NO) resulting from the iCIPT2 calculations were taken here as guidance. Those NOs with ONs ranging from 1.98 to 0.02 were carefully analyzed as candidates for active orbitals. However, to make the MR calculations feasible, the number of active orbitals was kept below 15 (by deleting those virtual orbitals of the smallest ONs). The resulting active spaces for the 24 radicals are documented in Table 4.3. The SA-CASSCF calculations can be initiated and supervised by the iCAS approach (imposed automatic selection of complete active spaces)293, where once the valence atomic orbitals (see Table S2 in the Supporting Information) are prepared, all remaining steps are automated. In particular, the subspace matching algorithm therein preserves the characters of the orbitals (doubly occupied, active, or virtual) throughout the self-consistent iterations. The CASSCF energies are documented in Table S3 in the Supporting Information.
[] Active spaces and electronic states by SA-CASSCF Radical Active Space Symmetry Irrep of Active MOa Electronic Stateb Doublet Quartet C3H5 (3e,9o) (2,2,1,4) (2,1,1,4) (0,1,1,0) BeF (9e,10o) (6,0,2,2) (2,0,1,1) - BeH (3e,11o) (5,0,3,3) (3,0,4,4) (0,0,1,1) BH2 (5e,10o) (5,0,3,2) (3,2,2,2) (0,1,0,0) CH (5e,9o) (5,0,2,2) (3,2,1,1) (0,1,1,1) CH3 (7e,11o) (6,0,3,2) (3,0,2,2) - CN (9e,8o) (4,0,2,2) (2,0,2,2) (2,2,0,0) CNO (15e,12o) (6,0,3,3) (3,2,3,3) (0,0,1,1) CON (15e,12o) (6,0,3,3) (1,1,2,2) (0,0,1,1) CO+ (9e,8o) (4,0,2,2) (2,0,1,1) (1,0,0,0) F2BO (19e,14o) (6,1,4,3) (2,1,2,1) - F2BS (19e,14o) (6,1,4,3) (1,0,2,1) (0,0,1,0) H2BO (11e,10o) (5,0,3,2) (2,0,3,2) (0,0,1,0) HCO (11e,10o) (8,2) (3,3) (1,1) HOC (11e,10o) (8,2) (3,3) (1,1) H2PO (13e,10o) (7,3) (4,3) (2,1) H2PS (13e,10o) (7,3) (3,2) (3,3) NCO (15e,12o) (6,0,3,3) (2,0,3,3) (0,0,1,1) NH2 (7e,10o) (5,0,3,2) (2,0,1,2) (0,0,0,1) CH2NO2 (13e,10o) (3,1,3,3) (2,2,2,2) (1,1,1,0) NO (11e,10o) (6,0,2,2) (2,0,2,2) (0,1,1,1) OH (7e,9o) (5,0,2,2) (1,1,1,1) (0,1,0,0) PH2 (7e,10o) (5,0,3,2) (2,1,2,2) (0,1,0,1) C2H3 (11e,12o) (10,2) (3,3) (1,1)
a
: (); : (). b: (); : ().
4.4 Vertical Excitation Energies
The SS-NEVPT2, MS-NEVPT2, SDSPT2, SDSCI, and ic-MRCISD calculations on the VEEs of the 110 doublet and 39 quartet states of the 24 radicals in QUEST#4145 were performed with the above setup of active spaces. The so-calculated VEEs are documented in Table LABEL:table-ene. Apart from the overall error analysis shown in Table 4.4, the internal consistency of the methods is also analyzed in Table 4.4. Scatter plots for the iCIPT2 and X-TDA/SS-NEVPT2/MS-NEVPT2/SDSPT2/SDSCI/ic-MRCISD VEEs are further depicted in Fig. 3.
It can be seen from Fig. 3 that the three variants of MRPT2 (i.e., SS-NEVPT2, MS-NEVPT2, and SDSPT2) have a strong linear correlation with iCIPT2, with R2 larger than 0.995, which is accompanied by a mean absolute deviation (MAD) of ca. 0.1 eV. As can be seen from Table 4.4, the VEEs calculated by the three MRPT2 are closely aligned for more than 80% of the excited states, with the differences in between being less than 0.05 eV. However, SS-NEVPT2 fails occasionally. For instance, it fails to reproduce the ordering of the (6.83 eV) and (6.71 eV) states of H2PO. Despite some overestimates, both MS-NEVPT2 and SDSPT2 reproduce the correct ordering for the two states (MS-NEVPT2: 6.61 and 7.04 eV; SDSPT2: 6.64 and 7.05 eV), as compared to the iCIPT2 results (6.20 and 6.81 eV). A similar situation occurs also to ( and under ) and ( and under ) states of CH3. These represent examples where dynamic correlation revises significantly the coefficients of the reference functions, as shown in Table 4.4. Here, it deserves to be emphasized that, although MS-NEVPT2 and SDSPT2 are hardly distinguishable for the systems considered here (see Table 4.4), SDSPT2 does outperform MS-NEVPT2 for situations with multiple near-degenerate states260.
Both SDSCI and ic-MRCISD have a perfect linear correlation with iCIPT2, with R2 being ca. 0.999, which is accompanied by a MAD of ca. 0.05 eV. SDSCI agrees with ic-MRCISD within 0.15 eV for all the excited states, given that the computational cost of SDSCI is merely that of one iteration of ic-MRCISD. More specifically, SDSCI agrees with ic-MRCISD within 0.05, 0.10, and 0.15 eV for 75%, 95%, and 100% of the excited states, respectively (cf. Table 4.4). In particular, this is not fortuitous but holds also for other systems147. However, it deserves to be pointed out that both SDSCI and ic-MRCISD perform worse than the three MRPT2 for the state of nitromethyl (CH2NO2), see Table LABEL:table-ene.
Compared to the above MR methods, X-TDA/BHandHLYP has a weaker linear correlation with iCIPT2, as reflected by a lower R2 of ca. 0.927. Moreover, the orderings of several states of BH2, CH3, H2PO, H2PS, and C2H3 were not reproduced correctly, as marked in Table LABEL:table-ene. Nonetheless, the accuracy of X-TDA for doublet and quartet excited states of radicals is comparable to that of TD-DFT for singlet and triplet excited states of closed-shell systems, as already noticed before200, 201.
[…]
As for the MRPT2 calculations, it is first noted that, although SS-NEVPT2 predicts correctly the relative ordering of the four states, it overestimates the excitation energies of the and states by 0.41 eV and 0.47 eV, respectively, but underestimates that of by 0.64 eV. To understand such failures of standard MRPT2, we analyzed the wavefunctions in detail. As can be seen from Table 4.4, the four states stem primarily from four transitions, two valence and two Rydberg transitions. For the state, which is predominantly a mixture of the two valence transitions, all MR methods yield similar results, with deviations from iCIPT2 by no more than 0.15 eV. In contrast, the wavefunctions of the other three states are more complex. At the CASSCF level, both and are dominated by Rydberg characters (>70%), while is more akin to a valence state yet mixed with 32% Rydberg character. The situation changes dramatically at the SDSCI and ic-MRCISD levels, where these states exhibit a strong mixing of all the four components, resulting in comparable Rydberg characters exceeding 50% for each state. Such revisions of the reference states cannot be captured by standard MRPT2 like SS-NEVPT2, but can be captured to some extents by their multi-state/quasi-degenerate variants. For instance, SDSPT2 reproduces essentially all the features of SDSCI/ic-MRCISD, whereas MS-NEVPT2 is much less successful, in the sense that its state contains only 25% Rydberg component, less than half of that of ic-MRCISD. Nevertheless, there is no significant difference in the overall accuracy of MS-NEVPT2 and SDSPT2: compared with iCIPT2, MS-NEVPT2 overestimates the excitation energies of the to states by 0.07, 0.12, and 0.07 eV, respectively, while SDSPT2 exhibits deviations of 0.22, 0.15, and -0.07 eV, respectively. It appears that MS-CASPT2 also performs well for these states298.
Radical | State | X-TDA | SS-NEVPT2 | MS-NEVPT2 | SDSPT2 | SDSCI | ic-MRCISD | iCIPT2a | 95% intervalb |
Allyl | -0.24 | 0.19 | 0.19 | 0.16 | -0.15 | -0.06 | 4.94 0.04 | 0.03 | |
(C3H5) | -0.28 | 0.24 | 0.24 | 0.21 | -0.11 | -0.03 | 5.46 0.06 | 0.03 | |
-0.36 | 0.18 | 0.18 | 0.16 | -0.17 | -0.08 | 5.61 0.01 | 0.03 | ||
0.07 | 0.16 | 0.15 | 0.10 | -0.13 | -0.05 | 3.40 0.02 | 0.03 | ||
-0.39 | 0.41 | 0.07 | 0.22 | 0.03 | 0.08 | 5.71 0.05 | 0.02 | ||
-0.60 | 0.47 | 0.12 | 0.15 | 0.07 | 0.07 | 6.53 0.01 | 0.03 | ||
-1.52 | -0.64 | 0.07 | -0.07 | -0.08 | -0.11 | 7.80 0.02 | 0.04 | ||
-0.31 | 0.22 | 0.22 | 0.19 | -0.25 | -0.12 | 5.98 0.02 | 0.02 | ||
-0.50 | 0.24 | 0.24 | 0.21 | -0.27 | -0.14 | 7.68 0.02 | 0.02 | ||
BeFc | -0.33 | 0.01 | 0.01 | 0.01 | -0.03 | -0.03 | 6.27 0.00 | 0.01 | |
0.05 | 0.02 | 0.02 | 0.04 | 0.06 | 0.03 | 4.15 0.00 | 0.00 | ||
BeHc | -0.37 | 0.01 | 0.01 | 0.01 | 0.01 | 0.00 | 5.52 0.00 | 0.00 | |
0.45 | 0.19 | 0.19 | 0.19 | 0.14 | 0.00 | 5.72 0.00 | 0.00 | ||
0.08 | 0.03 | 0.03 | 0.03 | 0.01 | 0.00 | 2.49 0.00 | 0.00 | ||
-0.76 | 0.00 | 0.00 | 0.01 | 0.01 | 0.00 | 6.46 0.00 | 0.00 | ||
0.44 | 0.24 | 0.23 | 0.23 | 0.16 | 0.01 | 7.34 0.00 | 0.00 | ||
0.12 | 0.19 | 0.20 | 0.21 | 0.11 | -0.03 | 7.80 0.00 | 0.00 | ||
0.24 | 0.03 | 0.03 | 0.03 | -0.01 | 0.00 | 5.88 0.00 | 0.00 | ||
BH2 | -0.24 | -0.03 | -0.02 | -0.02 | -0.01 | 0.01 | 5.72 0.00 | 0.00 | |
-0.35 | -0.04 | -0.03 | -0.03 | -0.02 | 0.01 | 6.75 0.00 | 0.00 | ||
0.17 | -0.03 | -0.02 | -0.02 | 0.00 | 0.01 | 6.41 0.00 | 0.00 | ||
0.08 | 0.06 | 0.06 | 0.06 | 0.04 | -0.03 | 6.95 0.00 | 0.00 | ||
0.19d | 0.05 | 0.05 | 0.06 | 0.04 | 0.00 | 6.38 0.00 | 0.00 | ||
-0.40d | -0.04 | -0.03 | -0.04 | -0.03 | 0.00 | 6.58 0.00 | 0.00 | ||
0.12 | 0.00 | 0.00 | 0.00 | 0.00 | 0.01 | 1.18 0.00 | 0.00 | ||
-0.37 | -0.06 | -0.05 | -0.05 | -0.02 | 0.01 | 6.87 0.00 | 0.00 | ||
0.19 | -0.03 | -0.02 | -0.02 | 0.00 | 0.01 | 5.35 0.00 | 0.00 | ||
CH | -1.05 | 0.06 | 0.05 | 0.05 | 0.02 | 0.00 | 3.98 0.00 | 0.00 | |
-0.10 | -0.02 | -0.02 | -0.02 | -0.01 | -0.01 | 6.49 0.00 | 0.00 | ||
0.47 | 0.03 | 0.03 | 0.03 | 0.01 | -0.01 | 2.91 0.00 | 0.00 | ||
0.10 | 0.02 | 0.02 | 0.03 | 0.01 | -0.01 | 3.29 0.00 | 0.00 | ||
0.11 | -0.03 | -0.03 | -0.04 | -0.01 | 0.01 | 0.72 0.00 | 0.00 | ||
-0.18 | 0.03 | 0.03 | 0.02 | 0.00 | 0.00 | 7.66 0.00 | 0.00 | ||
CH3 | -0.23 | -0.02 | -0.01 | -0.02 | -0.02 | 0.01 | 5.86 0.00 | 0.00 | |
0.04d | 0.17d | 0.07 | 0.09 | 0.11 | 0.04 | 6.96 0.00 | 0.00 | ||
-0.45d | -0.14d | -0.01 | -0.05 | -0.05 | -0.03 | 7.19 0.00 | 0.00 | ||
-0.26 | -0.06 | -0.03 | -0.03 | -0.01 | 0.00 | 7.65 0.00 | 0.00 | ||
CN | 0.36 | 0.02 | 0.04 | 0.05 | 0.02 | -0.01 | 3.22 0.00 | 0.00 | |
-0.46 | 0.04 | 0.05 | 0.05 | 0.00 | -0.01 | 1.34 0.00 | 0.00 | ||
0.84 | 0.11 | 0.12 | 0.14 | 0.07 | 0.01 | 7.88 0.00 | 0.00 | ||
-0.30 | 0.04 | 0.05 | 0.05 | 0.02 | 0.02 | 6.03 0.00 | 0.00 | ||
-0.70 | 0.07 | 0.08 | 0.09 | 0.03 | 0.00 | 7.12 0.00 | 0.00 | ||
-0.74 | 0.13 | 0.14 | 0.15 | 0.05 | 0.01 | 7.88 0.00 | 0.00 | ||
CNO | 1.49 | -0.08 | -0.08 | -0.08 | 0.05 | 0.01 | 1.62 0.01 | 0.00 | |
0.33 | 0.10 | 0.11 | 0.11 | 0.06 | 0.08 | 7.63 0.03 | 0.01 | ||
0.75 | 0.10 | 0.11 | 0.13 | 0.12 | 0.07 | 7.69 0.02 | 0.01 | ||
-0.43 | -0.19 | -0.18 | -0.16 | -0.18 | -0.18 | 7.82 0.02 | 0.01 | ||
0.31 | 0.04 | 0.05 | 0.05 | 0.02 | 0.02 | 5.48 0.01 | 0.00 | ||
0.00 | 0.09 | 0.10 | 0.11 | 0.05 | 0.06 | 6.18 0.03 | 0.01 | ||
0.08 | 0.07 | 0.08 | 0.09 | 0.04 | 0.06 | 5.70 0.01 | 0.00 | ||
CON | 2.42 | -0.04 | -0.04 | -0.04 | 0.03 | 0.00 | 3.80 0.02 | 0.01 | |
-0.47 | 0.04 | 0.04 | 0.07 | 0.07 | 0.05 | 7.15 0.02 | 0.01 | ||
-0.08 | 0.04 | 0.04 | 0.05 | 0.05 | 0.05 | 3.47 0.01 | 0.00 | ||
-0.20 | -0.03 | -0.03 | -0.01 | 0.01 | 0.02 | 2.74 0.00 | 0.00 | ||
CO+ | 0.83 | 0.05 | 0.05 | 0.07 | 0.01 | -0.01 | 5.80 0.00 | 0.00 | |
0.10 | 0.03 | 0.04 | 0.05 | -0.01 | 0.00 | 3.27 0.00 | 0.00 | ||
-0.27 | 0.06 | 0.07 | 0.08 | 0.02 | 0.01 | 7.27 0.00 | 0.00 | ||
F2BO | 0.15 | 0.04 | 0.04 | 0.04 | 0.01 | 0.01 | 2.77 0.01 | 0.04 | |
1.25 | 0.21 | 0.21 | 0.23 | 0.04 | 0.11 | 7.79 0.12 | 0.04 | ||
1.23 | 0.21 | 0.21 | 0.22 | -0.04 | 0.07 | 6.94 0.07 | 0.03 | ||
1.37 | 0.26 | 0.26 | 0.27 | -0.02 | 0.08 | 6.53 0.11 | 0.03 | ||
0.07 | -0.01 | -0.01 | -0.01 | -0.01 | 0.00 | 0.70 0.04 | 0.03 | ||
F2BS | 0.09 | 0.09 | 0.09 | 0.10 | -0.01 | 0.04 | 2.95 0.02 | 0.05 | |
0.09 | 0.08 | 0.08 | 0.10 | 0.16 | 0.12 | 7.02 0.05 | 0.04 | ||
-0.04 | 0.05 | 0.05 | 0.05 | -0.01 | -0.02 | 0.50 0.00 | 0.04 | ||
-0.03 | -0.04 | -0.04 | -0.03 | 0.08 | 0.05 | 6.39 0.03 | 0.05 | ||
H2BO | 0.02 | 0.06 | 0.06 | 0.07 | 0.01 | 0.00 | 3.51 0.00 | 0.01 | |
1.16 | 0.06 | 0.07 | 0.09 | 0.07 | 0.04 | 7.25 0.00 | 0.01 | ||
1.03 | 0.00 | 0.00 | 0.00 | 0.04 | 0.02 | 4.61 0.00 | 0.01 | ||
1.09 | 0.05 | 0.06 | 0.08 | 0.00 | 0.03 | 7.64 0.00 | 0.01 | ||
-0.21 | 0.04 | 0.04 | 0.05 | -0.02 | 0.00 | 2.17 0.00 | 0.01 | ||
0.98 | 0.08 | 0.08 | 0.11 | 0.07 | 0.05 | 6.13 0.01 | 0.01 | ||
0.73 | 0.07 | 0.07 | 0.09 | 0.00 | 0.05 | 6.86 0.00 | 0.01 | ||
HCO | 0.24 | 0.17 | 0.17 | 0.17 | -0.05 | 0.03 | 5.45 0.01 | 0.01 | |
0.07 | 0.17 | 0.18 | 0.18 | 0.07 | 0.08 | 6.15 0.01 | 0.01 | ||
0.17 | -0.01 | -0.02 | 0.00 | 0.04 | 0.02 | 2.09 0.00 | 0.01 | ||
-0.18 | 0.11 | 0.11 | 0.11 | -0.02 | 0.01 | 7.07 0.01 | 0.01 | ||
-0.28 | 0.08 | 0.10 | 0.11 | -0.01 | 0.00 | 7.73 0.01 | 0.01 | ||
-0.21 | 0.05 | 0.05 | 0.05 | -0.02 | 0.04 | 6.75 0.01 | 0.01 | ||
-0.21 | 0.04 | 0.05 | 0.05 | 0.00 | 0.05 | 6.37 0.01 | 0.01 | ||
HOC | 0.36 | 0.14 | 0.14 | 0.13 | -0.04 | 0.02 | 3.78 0.00 | 0.01 | |
0.11 | 0.13 | 0.13 | 0.14 | 0.09 | 0.04 | 5.63 0.01 | 0.01 | ||
0.01 | -0.01 | -0.01 | 0.00 | 0.02 | 0.00 | 0.92 0.00 | 0.01 | ||
0.20 | 0.07 | 0.07 | 0.10 | 0.06 | 0.03 | 6.13 0.01 | 0.01 | ||
0.28 | 0.15 | 0.15 | 0.17 | 0.08 | 0.06 | 7.55 0.01 | 0.01 | ||
0.24 | 0.18 | 0.18 | 0.17 | -0.02 | 0.05 | 7.73 0.01 | 0.01 | ||
0.05 | -0.08 | -0.08 | -0.08 | 0.00 | 0.02 | 3.84 0.01 | 0.01 | ||
H2PO | -0.05 | 0.18 | 0.24 | 0.25 | 0.08 | 0.04 | 4.21 0.01 | 0.03 | |
0.41 | 0.08 | 0.18 | 0.20 | 0.06 | 0.05 | 4.78 0.01 | 0.04 | ||
-0.03 | 0.14 | 0.21 | 0.23 | 0.02 | 0.03 | 5.66 0.02 | 0.02 | ||
-0.01 | 0.11 | 0.17 | 0.16 | -0.04 | 0.01 | 2.81 0.01 | 0.03 | ||
0.37 | 0.63d | 0.41 | 0.44 | 0.30 | 0.21 | 6.20 0.02 | 0.04 | ||
0.65 | -0.11d | 0.23 | 0.23 | -0.04 | 0.05 | 6.81 0.03 | 0.03 | ||
0.06 | 0.18 | 0.24 | 0.25 | 0.02 | 0.10 | 7.28 0.02 | 0.03 | ||
0.57 | 0.27 | 0.32 | 0.32 | 0.10 | 0.18 | 7.64 0.02 | 0.03 | ||
0.53 | 0.16 | 0.22 | 0.23 | 0.02 | 0.12 | 6.31 0.02 | 0.03 | ||
H2PS | 0.13 | 0.09 | 0.11 | 0.11 | 0.04 | 0.02 | 2.73 0.01 | 0.01 | |
0.12 | 0.19 | 0.20 | 0.21 | 0.02 | 0.02 | 4.57 0.02 | 0.01 | ||
-0.09 | 0.08 | 0.09 | 0.09 | -0.02 | 0.02 | 1.14 0.00 | 0.01 | ||
0.59 | 0.12 | 0.13 | 0.14 | 0.04 | 0.09 | 5.42 0.03 | 0.01 | ||
0.66 | 0.17 | 0.18 | 0.17 | 0.03 | 0.10 | 5.74 0.01 | 0.01 | ||
0.68d | 0.23 | 0.19 | 0.22 | 0.07 | 0.07 | 6.02 0.01 | 0.01 | ||
-0.76d | 0.34 | 0.40 | 0.40 | 0.15 | 0.17 | 6.76 0.02 | 0.01 | ||
1.09d | 0.21 | 0.19 | 0.20 | 0.10 | 0.10 | 5.08 0.01 | 0.02 | ||
-0.34d | 0.46 | 0.49 | 0.51 | 0.23 | 0.22 | 5.73 0.02 | 0.01 | ||
-0.05 | 0.14 | 0.16 | 0.16 | -0.02 | 0.06 | 6.72 0.01 | 0.01 | ||
NCO | 0.22 | -0.04 | -0.03 | -0.03 | 0.05 | 0.01 | 2.87 0.01 | 0.03 | |
1.22 | 0.07 | 0.08 | 0.08 | 0.03 | 0.03 | 7.03 0.02 | 0.02 | ||
0.88 | 0.03 | 0.05 | 0.05 | 0.02 | 0.02 | 4.72 0.01 | 0.02 | ||
0.04 | 0.10 | 0.11 | 0.12 | 0.07 | 0.05 | 7.43 0.03 | 0.02 | ||
-0.12 | 0.06 | 0.07 | 0.08 | 0.04 | 0.06 | 6.79 0.01 | 0.02 | ||
NH2 | 0.08 | 0.02 | 0.02 | 0.02 | 0.00 | 0.00 | 2.12 0.00 | 0.00 | |
0.04 | 0.06 | 0.06 | 0.05 | 0.00 | 0.00 | 7.69 0.00 | 0.00 | ||
0.01 | 0.06 | 0.06 | 0.07 | 0.01 | 0.01 | 6.48 0.00 | 0.00 | ||
0.02 | 0.03 | 0.03 | 0.02 | -0.02 | 0.00 | 7.75 0.00 | 0.00 | ||
0.04 | 0.04 | 0.04 | 0.03 | 0.00 | 0.01 | 7.29 0.00 | 0.00 | ||
Nitromethyl | 0.84 | -0.03 | 0.03 | -0.01 | 0.21 | 0.15 | 2.50 0.01 | 0.03 | |
(CH2NO2) | 0.35 | -0.05 | 0.01 | -0.01 | 0.06 | 0.07 | 5.20 0.01 | 0.04 | |
1.13 | -0.05 | -0.02 | -0.04 | 0.10 | 0.08 | 2.30 0.00 | 0.03 | ||
0.64 | 0.03 | 0.13 | 0.15 | 0.11 | 0.15 | 6.70 0.07 | 0.03 | ||
1.02 | -0.03 | 0.03 | 0.01 | 0.13 | 0.11 | 2.00 0.02 | 0.04 | ||
0.33 | -0.02 | 0.04 | 0.02 | -0.01 | 0.03 | 4.66 0.02 | 0.04 | ||
0.21 | -0.05 | 0.08 | 0.13 | 0.42 | 0.30 | 5.31 0.00 | 0.03 | ||
0.04 | 0.00 | 0.06 | 0.04 | 0.06 | 0.09 | 4.94 0.03 | 0.02 | ||
-0.52 | -0.05 | 0.01 | -0.02 | -0.05 | 0.00 | 4.30 0.02 | 0.03 | ||
0.10 | 0.01 | 0.07 | 0.05 | 0.00 | 0.05 | 4.46 0.03 | 0.04 | ||
NO | 0.46 | 0.19 | 0.19 | 0.19 | -0.09 | -0.01 | 6.11 0.01 | 0.01 | |
0.28 | 0.17 | 0.17 | 0.16 | -0.11 | -0.02 | 7.28 0.01 | 0.00 | ||
-0.45 | -0.04 | -0.04 | -0.04 | -0.15 | -0.13 | 7.79 0.00 | 0.01 | ||
-0.21 | -0.04 | -0.04 | -0.02 | 0.05 | 0.02 | 6.36 0.00 | 0.01 | ||
-0.45 | 0.06 | 0.06 | 0.05 | -0.03 | 0.02 | 6.78 0.00 | 0.01 | ||
OH | 0.04 | 0.01 | 0.01 | 0.01 | 0.01 | 0.00 | 4.11 0.00 | 0.00 | |
0.09 | 0.01 | 0.01 | 0.01 | -0.01 | -0.01 | 8.03 0.00 | 0.00 | ||
0.09 | 0.02 | 0.02 | 0.02 | 0.00 | -0.01 | 7.50 0.00 | 0.00 | ||
PH2 | 0.05 | 0.07 | 0.07 | 0.07 | 0.04 | -0.01 | 2.76 0.00 | 0.00 | |
-0.15 | 0.05 | 0.05 | 0.04 | 0.00 | 0.02 | 5.90 0.00 | 0.00 | ||
0.24 | 0.21 | 0.21 | 0.22 | 0.10 | 0.02 | 7.10 0.00 | 0.00 | ||
0.08 | 0.10 | 0.10 | 0.10 | 0.03 | -0.02 | 5.17 0.00 | 0.00 | ||
-0.07 | 0.08 | 0.09 | 0.09 | 0.02 | -0.02 | 5.75 0.00 | 0.00 | ||
-0.14 | 0.06 | 0.06 | 0.06 | 0.01 | 0.01 | 7.47 0.00 | 0.00 | ||
0.09 | 0.04 | 0.04 | 0.04 | 0.00 | 0.03 | 6.16 0.00 | 0.00 | ||
-0.03 | 0.13 | 0.13 | 0.13 | 0.06 | 0.05 | 6.95 0.00 | 0.00 | ||
Vinyl | 0.75d | 0.10 | 0.09 | 0.11 | 0.02 | 0.02 | 5.60 0.00 | 0.02 | |
(C2H3) | -0.80d | 0.13 | 0.17 | 0.15 | -0.06 | 0.01 | 6.19 0.01 | 0.01 | |
-0.10 | 0.04 | 0.05 | 0.05 | 0.01 | -0.03 | 3.28 0.00 | 0.01 | ||
-0.03 | 0.05 | 0.06 | 0.09 | 0.08 | 0.02 | 4.70 0.00 | 0.01 | ||
-0.42 | 0.14 | 0.16 | 0.15 | -0.07 | 0.02 | 7.54 0.01 | 0.01 | ||
-0.15 | 0.05 | 0.06 | 0.05 | -0.03 | 0.02 | 4.54 0.01 | 0.01 | ||
-0.38 | 0.16 | 0.17 | 0.15 | -0.08 | 0.03 | 7.40 0.01 | 0.01 |
a
means extrapolation distance .
b
Sum of the 95% confidence intervals of ground and excited states.
c
The AVTZ basis set of Be varies across different program packages.
d
Incorrect ordering.
[] Statistical analysis of the errors (in eV) of X-TDA, SS-NEVPT2, MS-NEVPT, SDSPT2, SDSCI, and ic-MRCISD relative to iCIPT2 for the vertical excitation energies of 110 doublet and 39 quartet states∗ Type X-TDA SS-NEVPT2 MS-NEVPT2 SDSPT2 SDSCI ic-MRCISD Doublets MD 0.17 0.07 0.08 0.08 0.02 0.02 (110 states) MAD 0.40 0.10 0.09 0.10 0.06 0.04 SD 0.58 0.15 0.12 0.12 0.08 0.06 MAX 2.42 0.63 0.41 0.44 0.42 0.30 Quartets MD -0.03 0.09 0.11 0.11 0.01 0.04 (39 states) MAD 0.31 0.11 0.12 0.12 0.05 0.06 SD 0.41 0.15 0.16 0.16 0.08 0.08 MAX 1.09 0.46 0.49 0.51 0.23 0.22 Overall MD 0.12 0.07 0.09 0.09 0.02 0.03 (149 states) MAD 0.38 0.10 0.10 0.10 0.05 0.04 SD 0.54 0.15 0.13 0.13 0.08 0.07 MAX 2.42 0.63 0.49 0.51 0.42 0.30
MD: mean deviation; MAD: mean absolute deviation; SD: standard deviation; MAX: maximum deviation.
[] Internal consistency (percentage number of states within a deviation range in absolute value) across SS-NEVPT2, MS-NEVPT2, SDSPT2, SDSCI, ic-MRCISD, and iCIPT2 Deviation Range (eV) SS-NEVPT2 vs MS-NEVPT2 SS-NEVPT2 vs SDSPT2 MS-NEVPT2 vs SDSPT2 MS-NEVPT2 vs SDSCI SDSPT2 vs SDSCI SDSCI vs ic-MRCISD ic-MRCISD vs iCIPT2 0.05 84 % 89 % 99 % 50 % 44 % 75 % 74% 0.10 95 % 95 % 99 % 72 % 68 % 95 % 89% 0.15 97 % 96 % 100 % 82 % 81 % 100 % 96%
[] CASSCF, MS-NEVPT2, and SDSPT2 wavefunctions of excited states of H2PO and CH3 H2PO Ground state: (10a’)1(3a”)2 (3a”)2 (11a’)0 (10a’)1 (4a”)1 72% 15% 18% 62% 3% 75% 87% 3% 7% 73% 84% 5% CH3 Ground state: (3a1)2(1b2)2(1b1)1 (1b1)1 (5a1)0 (3a1)2 (1b1)1 96% 0% 0% 95% 20% 76% 75% 20% 42% 54% 54% 41% (1b1)1 (2b2)0 (1b2)2 (1b1)1 96% 0% 0% 95% 20% 76% 75% 20% 48% 48% 48% 47%
[] Vertical excitation energies (in eV) of C3H5 Typea SS-NEVPT2 MS-NEVPT2 SDSPT2 SDSCI ic-MRCISD iCIPT2b 95% intervalc AVTZ AVTZ AVTZ AVTZ AVTZ AVTZ Val. 3.57 3.56 3.50 3.28 3.36 3.40 0.02 0.03 Ryd. 5.13 5.13 5.10 4.79 4.87 4.94 0.04 0.03 Ryd. 5.70 5.70 5.68 5.35 5.44 5.46 0.06 0.03 Ryd. 5.79 5.79 5.77 5.43 5.53 5.61 0.01 0.03 Ryd. 6.12 5.78 5.93 5.74 5.79 5.71 0.05 0.02 Val. 6.20 6.20 6.17 5.73 5.86 5.98 0.02 0.02 Ryd. 7.00 6.65 6.68 6.60 6.60 6.53 0.01 0.03 Ryd. 7.92 7.92 7.89 7.41 7.54 7.68 0.02 0.02 Ryd. 7.17 7.87 7.73 7.72 7.69 7.80 0.02 0.04 Typea CI2d ic-MRCISDe MS-CASPT2f ROCC3h mcf effecti iCIPT2b 95% intervalc DZP DZP ANO-L AVTZ of UCCSD AVTZ + 2s1p on C + 1s1p on C + 1s1p1d on C + mcf: 2s2p2d Val. 3.36 3.32 3.32 3.44 0.03 3.40 0.02 0.03 Ryd. 4.70 4.68 5.11 4.94 -0.01 4.94 0.04 0.03 Ryd. 5.25 5.25 5.65 5.49 -0.02 5.46 0.06 0.03 Ryd. 5.31 5.29 5.76 5.59 0.09 5.61 0.01 0.03 Ryd. 5.50 5.46 5.73 5.65 -0.09 5.71 0.05 0.02 Val. - - 5.89g 6.15 0.04 5.98 0.02 0.02 Ryd. - - 6.36 6.20 -0.49 6.53 0.01 0.03 Ryd. - - - - - 7.68 0.02 0.02 Ryd. - - 6.90 6.79 -1.44 7.80 0.02 0.04
a
Val.: valence excitation; Ryd.: Rydberg excitation.
b
means extrapolation distance .
c
Sum of the 95% confidence intervals of ground and excited states.
d
Ref. 300; second-order MRCI with (3e,7o).
e
Ref. 301; ic-MRCISD with (3e,7o).
f
Ref. 298; ANO-L: C(14s9p4d)/H(8s4p) C[4s3p2d]/H[3s2p]; (3e,7o) for the doublets, whereas (3e,3o) for .
g
Ref. 298; CASPT2 with (3e,3o).
h
Ref. 299; mcf: molecule-centered basis functions.
i
Ref. 299; E(UCCSD/AVTZ + mcf) E(UCCSD/AVTZ).
[] CASSCF, MS-NEVPT2, SDSPT2, SDSCI, and ic-MRCISD wavefunctions of excited states of C3H5 C3H5 Ground state: (6a1)2(1a2)1(4b2)2(1b1)2 (1a2)1 (2b1)0 (1b1)2 (1a2)1 (1a2)1 (3b1)0 (1a2)1 (4b1)0 Rydberg Valence Valence Rydberg Rydberg component 41% 49% 1% 0% 1% 10% 5% 51% 28% 79% 8% 11% 41% 33% 74% 29% 25% 0% 32% 32% 43% 48% 0% 0% 0% 37% 30% 11% 14% 25% 6% 10% 71% 3% 74% 2% 3% 10% 76% 86% 41% 49% 0% 0% 0% 21% 13% 34% 26% 60% 14% 17% 55% 4% 59% 11% 12% 3% 63% 66% 40% 51% 0% 0% 0% 18% 9% 39% 28% 67% 14% 15% 52% 11% 63% 17% 16% 2% 54% 56% 36% 45% 0% 0% 0% 18% 11% 32% 23% 55% 13% 15% 48% 5% 53% 11% 11% 3% 55% 58%
5 Conclusions
A comprehensive dataset, QUEST#4X, has been introduced for the purpose of calibrating multireference (MR) methods. It encompasses 110 doublet and 39 quartet excited states for the 24 radicals in the dataset QUEST#4 (which has only 51 doublet states). The iCIPT2/AVTZ vertical excitation energies reported here serve as a benchmark for any other MR methods. It is of interest to note that the MR methods considered here, i.e., NEVPT2, SDSPT2, SDSCI, and ic-MRCISD, share the same function space and differ only in the determination of the expansion coefficients. Because of this, SDSCI can be taken as the initial step of ic-MRCISD and can even yield SDSPT2 and MS-NEVPT2 results automatically. Moreover, SDSPT2 also yields MS-NEVPT2 results automatically. More interestingly, the degree of agreement between SDSPT2 and MS-NEVPT2, the magnitude of deviations of SDSPT2/MS-NEVPT2 from SDSCI, as well as the iterative increments going from SDSCI to ic-MRCISD are clear indicators for the internal consistency between the methods and hence their accuracy. A good internal consistency can only be achieved by choosing appropriate active spaces for all target states, as done here. Notwithstanding this, more challenging datasets for the excited states of open-shell transition metals and -elements remain to be established. Work along this direction is being carried out at our laboratory.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 22373057) and Mount Tai Scholar Climbing Project of Shandong Province.
Supporting Information
Information for the radicals under concern, proper choice of zeroth-order Hamiltonian for Rydberg states, valence atomic orbitals, and CASSCF energies.
Conflicts of interest
There are no conflicts to declare.